A stone is thrown vertically upwards with an initial velocity u from the top of a tower,reaches the ground with a velocity of 3u.The height of the tower is???

Grade:11

3 Answers

IHate Ppts
6 years ago
When it reaches maximum height and falls down till the point of projection, the velocity at that instant and projection will be same. Using that logic, we will solve it. 
Firstly, we use V=U+AT to find time.
3u=u+10t, t=u/5
Substitute this time in displacement formula.
S=u(u/5)+1/2 * 10(u/5)2
s = 10u2/25
harshal pawar
6 years ago
It is solved by the equation v2=u2+2as 
 
And the answer 4u^2/g..
Cheers.
Dt
5 years ago
H be height v =velocity g= gravitational force or acceleration 
H=v2-u2/2g
H=9u2-u2/2g
H=8u2/2g
H=4u2/g is the height of the tower 
These answer is 100% accurate